The following offers information on the demand and cost structure for a monopolist. Determine the total...

  1. The following offers information on the demand and cost structure for a monopolist.
    1. Determine the total revenue and marginal revenue for the firm and fill in the table.
    2. Using the information, you just found, determine the profit maximizing price and quantity for the monopolist. Explain your answer.
    3. At the profit maximizing output, what is the profit/loss of the monopolist? Show your work.
    4. Compute the deadweight loss that results from the lack of competition in this market. Show your work.
    5. If the government were to regulate this monopolist and keep the price so that normal profits were zero, what price should the government set for the monopolist? Explain.
    6. If the government were to charge a price equal to marginal cost at the profit maximizing output, what would happen to the monopolist’s profit? Explain.

Quantity

Price per Unit

Total Cost

Marginal Cost

Total Revenue

Marginal Revenue

10

$10

$20

20

$8

$50

30

$6

$65

40

$4

$90

50

$2

$120
Answer #1

(a)

  • TR = P x Q
  • MR = Change in TR / Change in Q
  • MC = Change in TC / Change in Q
Q P TC MC TR MR
10 10 20 2 100 10
20 8 50 3 160 6
30 6 65 1.5 180 2
40 4 90 2.5 160 -2
50 2 120 3 100 -6

(b)

Profit is maximized when MR = MC, but this equality doesn't hold in this case.

When Q = 30, MR = 2 and MC = 1.5, so MR > MC.

When Q = 40, MR = - 2 and MC = 2.5, so MR < MC.

So profit is maximized when Q = 30 and P = 6.

(c)

When Q = 30,

Profit = TR - TC = 180 - 65 = 115

(d)

In efficient outcome, P = MC, but this equality doesn't hold in this case.

When Q = 40, P = 4 and MC = 2.5, so P > MC.

When Q = 50, P = 2 and MC = 3, so P < MC.

So efficiency is maximized when Q = 40 and P = 4.

Deadweight loss = (1/2) x Change in P x Change in Q

= (1/2) x (6 - 4) x (40 - 30) = (1/2) x 2 x 10 = 10

NOTE: As HOMEWORKLIB Answering Policy, 1st 4 parts are answered.

.